The mean height of a woman in a country (ages 20-29) is 64.4 inches. A random sample of 70 women in this age group is selected. What is the probability that the mean height for the sample is greater than 65 inches? Assume the mean is 2.88.
Answer by stanbon(75887)

The mean height of a woman in a country (ages 20-29) is 64.4 inches. A random sample of 70 women in this age group is selected. What is the probability that the mean height for the sample is greater than 65 inches? Assume the standard
deviation is 2.88.
-------
z(65) = (65-64.4)/[2.88/sqrt(70)] = 1.7430
P(x > 65) = P(z > 1.7430) = normalcdf(1.7430,100) = 0.0407
-------------
Cheers,
Stan H.
------------